# Isotope shift analysis with the $4f^{14}6s^{2}~{}^{1}S_{0}-4f^{13}5d6s^{2}(J=2)$ transition in ytterbium

### v1

@@ III Theoretical calculation of the electronic structure
The electronic structure of Yb is numerically calculated with configuration interaction implemented in ambit [21]. For Yb, up to $n=19$ states for $s$ and $p$ orbitals and $n=14$ states for $d$ and $f$ orbitals are included in the calculation. For ${}^{+}$, up to $n=16$ states are incorporated for all $s$, $p$, $d$, and $f$ orbitals. Two electron excitations from states relevant to this analysis and some other excited states and one hole excitations from the $4f$ orbital are allowed. For Yb, a hole excitation from the $5s$ and $5p$ orbitals are also allowed. $F_{\alpha}$, $K_{\alpha}$, and $D_{\alpha}$ are calculated by adjusting the initial conditions to observe the resulting energy shifts of each excited state.
In the following discussion, the results for Yb are mainly used, and the quantities for ${}^{+}$ are calculated to test consistency with previous reports. The calculations in this work are reasonably consistent with previous reports [7, 1, 6, 22] and experimental values [23], except for $K_{\alpha}$ for ${}^{+}$ ions and $K_{\beta\alpha}$, as shown in Table 6 in Appendix.
@@ IV.1 Hyperfine Structure
From the absolute frequencies listed in Table 2, hyperfine constants are calculated. The A constant for ${}^{171}$b is reported in Ref. [19]. For ${}^{173}$b, hyperfine levels are characterized by the A, B, and C hyperfine constants $A_{h}$, $B_{h}$, and $C_{h}$ corresponding to magnetic dipole, electric quadrupole, and magnetic otcupole components (see Appendix). The best fits of the absolute frequencies with and without the octupole term are summarized in Table 3. With the octupole term, the fit quality characterized by $\chi^{2}/{\rm ndf}$ improves. However, the obtained octupole moment has only a significance $1.62\sigma$, and the shift of the dipole and quadrupole constants induced by the introduction of the octupole term is within the standard deviation. Further improvement in the precision of the spectroscopy is desired to determine whether the nucleus has a finite octupole moment. In the following analysis, to be conservative, $A_{h}$ and $B_{h}$ obtained by the fit without the octupole term are used.
Table 3: Hyperfine constants for fermionic Yb isotopes in the unit of kHz: the uncertainties include the inflation by $\sqrt{\chi^{2}/{\rm ndf}}$.
without $C^{173}_{h}$ with $C^{173}_{h}$ Ref.
${}^{171}$b $A^{171}_{h}$ 1 123 273(13) [19]
${}^{173}$b $A^{173}_{h}$309 446.1(7.0) -309 449.1(5.5) this work
$B^{173}_{h}$1 700 631(78) -1 700 652(60)
@@
$\chi^{2}/{\rm ndf}$ 12.05 6.711
The ratio of $A_{h}$ between different isotopes equals the ratio of the nuclear $g$ factors $g^{A}$. The deviation $\Delta_{A,A^{\prime}}$ defined by the following equation is referred to as hyperfine anomaly [24]:
$\Delta_{173,171}=\frac{A_{h}^{173}/A_{h}^{171}}{g^{171}/g^{173}}-1=\frac{A_{h}^{173}/A_{h}^{171}}{(\mu_{n}^{171}/I^{171})/(\mu_{n}^{173}/I^{173})}-1,$ (5)
where $\mu_{n}^{A}$ and $I^{A}$ are the nuclear magnetic moment and nuclear spin, and $(\mu_{n}^{171}/I^{171})/(\mu_{n}^{173}/I^{173})=-3.6305(1)$ [25, 26, 27]. The hyperfine anomaly for the 431 nm and other transitions are summarized in Table 4. $\Delta_{173,171}$ for the $4f^{13}5d6s^{2}(J=2)$ state corresponding to the 431 nm transition is finite but particularly small compared to other states in Yb, as well as anomalies in most alkali atoms [24]. Also, it is worth noting that the singlet states tend to have large anomaly.
Table 4: Hyperfine anomalies $\Delta_{171,173}$ and electric quadrupole moments $Q$ for different states in Yb: references profvide values for the A and B constants.
[…]
$6s6p~{}^{1}P_{1}$ 13.90(56) 4664.6(2.5) [28]
$4f^{13}5d6s^{2}(J=2)$0.156(55) 404.634(27) this work
$6s6p~{}^{3}P_{2}$8.96(53) -133.066(56) [29]
$6s6p~{}^{3}P_{1}$3.856(70) 333.574(35) [30]
$6s7s~{}^{3}S_{1}$3.50(34) 0.047(233) [29]
$6s5d~{}^{1}D_{2}$13.0(6.7) 77.0(5.0) [31]
In the nonrelativistic limit, $A_{h}$ and $B_{h}$ are written in the following form [24].
$A_{h}$ $=$ $-\frac{1}{h}\frac{16\pi}{3}\frac{\mu_{0}}{4\pi}\mu_{B}\mu_{N}\frac{2L(L+1)}{J(J+1)}g^{A}\langle r^{-3}\rangle_{nl}$ (6)
$B_{h}$ $=$ $\frac{1}{h}\frac{e^{2}}{4\pi\epsilon_{0}}\frac{2J-1}{2J+2}\langle r^{-3}\rangle_{nl}Q$ (7)
Here, $h$ is the Planck constant, $\mu_{0}$ and $\epsilon_{0}$ are the magnetic and electric constants, $L$ and $J$ are the electronic orbital and total angular momentum, $e$ is the elementary charge, $\mu_{B}$ and $\mu_{N}$ are the Bohr and nulcear magnetons, and $\langle r^{-3}\rangle_{nl}$ is the average over the wavefunction of the electronic state $(n,l)$. The nuclear quadrupole moment $Q$ can then be calculated as
$Q=-\frac{\epsilon_{0}\mu_{0}}{e^{2}}\mu_{B}\mu_{N}g^{A}\frac{4L(L+1)}{J(2J-1)}\frac{B_{h}}{A_{h}}$ (8)
[…]
@@ IV.2 Nuclear charge radius
With theoretically calculated $F_{\alpha}$ and $K_{\alpha}$, $\langle r^{2}\rangle^{AA^{\prime}}$ is estimated using Eq. (1). The relative uncertainty of $\langle r^{2}\rangle^{AA^{\prime}}$ caused by the uncertainty of the spectroscopy is $\sim 10^{-5}$ (see Table 7 in Appendix), and the uncertainty for the obtained $\langle r^{2}\rangle^{AA^{\prime}}$ mainly originates from that of the theoretical calculation of $F_{\alpha}$ and $K_{\alpha}$, whose amount is difficult to estimate. To determine the uncertainty for $\langle r^{2}\rangle^{AA^{\prime}}$, $\langle r^{2}\rangle^{AA^{\prime}}_{\alpha}$ is calculated with isotope shift data for $\alpha=399$ [28], $431$, $556$ [30], and $578$ [18]. The average of them is regarded as the most likely $\langle r^{2}\rangle^{AA^{\prime}}$, and standard deviation of them is regarded as the uncertainty.
[…]
Figure 2: $\langle r^{2}\rangle^{AA^{\prime}}$ calculated from isotope shifts with $A^{\prime}=176$: red circles show the values obtained from the absolute frequencies of the 431 nm transition, and black triangles show the average over the 399, 431, 556, and 578 nm transitions. The inset shows $\Delta\langle r^{2}\rangle^{(3)}$.
From these data, three point odd-even staggering $\Delta\langle r^{2}\rangle^{(3)}$ is calculated as
@@
This is plotted in the inset of Fig. 2. Compared to the odd-even staggering observed in other atoms [40, 41, 42, 43], this effect is relatively large for stable isotopes. This could be attributed to the fact that both the number of protons and neutrons are far from the magic numbers (50 and 82).
@@ IV.3 King Plot
Figure 3: Normalized King plot for the 431 nm transition: the plot is normalized with the 578 nm transition. Texts near each data point show the pair of isotopes that generated the data point. The blue line shows the linear fit of the black data points. See Table 6 in Appendix for the fit parameters. The bottom half displays the residual of each data point from the fitted line. The residuals for the bosonic isotopes (black points) are magnified by a factor of 10.
The normalized King plot between the 431 nm transition and the 578 nm transition following Eq. 3 is shown in Fig. 3. In addition to the data points for bosonic isotopes, data points for fermionic isotopes are plotted. The representative transition frequency for a fermionic isotope is calculated as the center of gravity frequency of the hyperfine levels, and the isotope shift is calculated as the shift from the transition frequency for ${}^{174}$b, so that the data points lie between those of bosonic isotopes and have smaller uncertainties. First, to assess the level of nonlinearity of the King plot, linear fit for the bosonic isotopes is performed by Eq. 3 with only the first two terms on the right hand side. The plot exhibits slight nonlinearity beyond the standard deviation for some data points, as characterized by $\chi^{2}/{\rm ndf}=1.264$. However, as the probability for this level of deviation from $\chi^{2}/{\rm ndf}=1$ is 28.24%, the data points are linear within 95% confidence level.
@@
Although the amount of the nonlinearity is consistent with zero, the residual $d^{A}$ for isotope $A$ can be decomposed onto the two dimensional space made by $\zeta_{\pm}=d^{168}-d^{170}\pm(d^{172}-d^{174})$, as introduced in Ref. [7]. The nonlinearity of the 431 nm transition shown in Fig. 4 is consistent with the one-parameter fit for the 361, 411, 436 and 467 nm transitions within one standard deviation. This finding supports the argument made in Ref. [1] that the major part of the nonlinearities in the King plots for Yb originates from several nuclear effects. This consistency observed between the previous results, as well as the almost linear behavior on the King plot, ensures that the measurements are not significantly affected by any systematic shifts beyond the obtained uncertainties.
Table 5: Residual $\delta\nu^{A}$ from the linear fit of the King plot for fermionic isotopes: the residual on the plot is multiplied by $\nu^{AA^{\prime}}_{578}$ and displayed in the unit of kHz.
$\alpha$ $\delta\nu^{171}$ (kHz) $\delta\nu^{173}$ (kHz)
@@
$d^{A}$ for the fermionic isotopes is significantly larger than the uncertainties on it, as shown in the bottom half of Fig. 3. To investigate the origin of this offset, the normalized King plots are made for the 399 and 556 nm transitions as well, and $\delta\nu^{A}=d^{A}\nu_{578}^{AA^{\prime}}$ is extracted, as shown in Table 5. On one hand, it is at most $\sim 1$ MHz and thus not significant in older experiments with large uncertainties [44, 45, 46]. On the other hand, with an accuracy of $\sim 10$ kHz, this is significant. $\delta\nu^{171}$, $\delta\nu^{173}$, and $\delta\nu^{173}/\delta\nu^{171}$ are plotted against $Q$ in Table 4, $\Delta_{171,173}/10^{-3}$, $A_{h}^{171}$, $B_{h}^{173}$ , $B_{h}^{173}/A_{h}^{173}$, and $F_{\alpha}$ (See Figs. 8, 9, and 8 in Appendix). Quantities with largest correlation are $B^{173}$ for $\delta\nu_{173}$ and $F_{\alpha}$ for $\delta\nu_{171}$, respectively. $\delta\nu^{173}/\delta\nu^{173}$ does not look to have any good correlation against other quantities. However, it is not evident enough to state these quantities are related, because the $\chi^{2}/{\rm ndf}$ for the linear fit are 7.155 and 5.492, respectively, and both fitted lines do not cross the origin on the two-dimensional plot. One reason why the analysis is inconclusive is because $\delta\nu_{171}$ for the 399 nm transition has a large relative uncertainty. More precise spectroscopy, not only limited to the 399, 431, and 556 nm transitions but also for other narrow-linewidth transitions, is desired to figure out the source of the large $d^{A}$ for fermionic isotopes.

### v2

@@ III Theoretical calculation of the electronic structure
The electronic structure of Yb is numerically calculated with configuration interaction implemented in ambit [27]. For Yb, up to $n=19$ states for $s$ and $p$ orbitals and $n=14$ states for $d$ and $f$ orbitals are included in the calculation. For Yb+, up to $n=16$ states are incorporated for all $s$, $p$, $d$, and $f$ orbitals. Two electron excitations from states relevant to this analysis and some other excited states and one hole excitation from the $4f$ orbital is allowed. For Yb, a hole excitation from the $5s$ and $5p$ orbitals is also allowed. $F_{\alpha}$, $K_{\alpha}$, and $D_{\alpha}$ are calculated by adjusting the initial conditions to observe the resulting energy shifts of each excited state.
The results of our calculation are shown in Table 3, together with experimental measurements and results in other reports. Given the situation where sometimes $>50$% disagreements are observed between other reports and experimental measurements (see $\nu_{467}$ in Table 3), most of our calculations agree well with experimental values and other reports, within $<10$% relative differences. As for the 431-nm transition, $\nu_{431}$ and $F_{431}/F_{578}$ have experimental measurements. The relative offsets of the calculation from experiments for these two quantities are 17 and 13%, respectively. The obtained mass shift has large differences to other calculations. However, previous reports on theoretical calculation also have large discrepancy between each other with very large relative uncertainties in some cases. Given this, it might be necessary to conservatively assume ±$\pm 20$% relative uncertainty, but our calculation is overall reasonably compatible with previous reports.
Table 3: Result of the theoretical calculation of the electronic structure in Yb and Yb+: the units of $\nu_{\alpha}$, $F_{\alpha}$, $K_{\alpha}$, $D_{\alpha}$, $K_{\beta,\alpha}$, and $D_{\beta,\alpha}$ are THz, GHz/fm2, GHz u, $10^{3}$ THz, GHz u, and $10^{3}$ THz, respectively. $D_{\alpha}$ values are at $m_{\phi}=1$ eV. Experimental values for $\nu_{\alpha}$ are cited from Ref. [28], and the exact value differs by a few GHz according to isotope shifts. Experimental values for $F_{\beta}/F_{\alpha}$ and $K_{\beta\alpha}$ are obtained by the linear fit of King plots for the even-mass isotopes (see Section IV.3). In Refs. [29, 1, 2], the convention of the mass shift is different and a negative sign is added to compensate it.
this work ambit [3] grasp2018 [3] grasp2018 [30] Ref. [1] Ref. [2] Other reports Experiment
$\nu_{361}$ 875.19 819.47 829.76
$\nu_{399}$ 744.82 751.53
$\nu_{411}$ 691.88 707.00 808.11 729.48
$\nu_{431}$ 810.67 695.17
$\nu_{436}$ 649.86 679.86 770.13 688.35
$\nu_{467}$ 727.06 1051.44 580.12 642.12
$\nu_{556}$ 524.06 543.18 539.39
$\nu_{578}$ 500.53 522.78 458.36 522.68 518.29
$F_{361}$11.988 -13.528 -14.437 [19]
$F_{399}$7.020
$F_{411}$15.534 -14.715 -15.852 -17.604 [19]
$F_{431}$ 17.771
$F_{436}$15.770 -14.968 -16.094 -18.003 [19]
$F_{467}$ 35.260 36.218 41.892
$F_{556}$9.893 -10.951(21) -11.5 -9.3(2.1) -12.3(2) [29]
$F_{578}$9.692 -9.719 -9.1508 -10.848(21)
$K_{399}$ 412
$K_{411}$308 -752 -1678.2
$K_{431}$ 10172
$K_{436}$158 -661 -1638.5
$K_{467}$ 10167 12001 3127.6
$K_{556}$532 -280(72) -4581(1531) -580(3820) -1500(500) [29]
$K_{578}$527 -288(75) -655 [8]
$D_{411}$ 43.37 43.158 44.145 41.235 [19]
$D_{431}$194.34
$D_{436}$ 49.74 48.634 48.419 48.795 [19]
$D_{467}$295.06 -352.38 -730.4
$D_{578}$39.49 -42.855 -55.729
$F_{399}/F_{578}$ 0.7243 0.4453(12)
$F_{431}/F_{578}$1.8336 -1.62301(8)
$F_{556}/F_{578}$ 1.0207 1.0095(28) 1.00594(26)
$K_{399,578}$ 793.7 1332(17)
$K_{431,578}$ 9205.7 6430.9(1.4)
$K_{556,578}$ 5.929 10.7(104.5) 33.2(4.0)
$D_{431,578}$266.7
@@ IV.1 Hyperfine Structure
From the absolute frequencies listed in Table 2, hyperfine constants are calculated. The hyperfine A constant $A_{h}^{171}$ for 171Yb is reported in Ref. [22]. For 173Yb, first-order hyperfine interaction is characterized by the A, B, and C hyperfine constants $A_{h}^{173}$, $B_{h}^{173}$, and $C_{h}^{173}$ corresponding to the magnetic dipole, electric quadrupole, and magnetic octupole components (see Appendix B). The best fits of the absolute frequencies with and without the octupole term are summarized in Table 4. With the octupole term, the fit quality characterized by $\chi^{2}/{\rm ndf}$ (ndf: number of degrees of freedom) improves but is still somewhat poor with $\chi^{2}/{\rm ndf}=6.711$. The obtained octupole moment $C^{173}_{h}=7.6(4.7)$ kHz, which includes the inflation of the uncertainty with $\sqrt{\chi^{2}/{\rm ndf}}$, has only a significance $1.62\sigma$, and the shift of the dipole and quadrupole constants induced by the introduction of the octupole term is within the standard deviation after inflating it by $\sqrt{\chi^{2}/{\rm ndf}}$.
The effect of the second-order hyperfine interaction $W_{F}^{(2)}$ can also be significant in Yb (see Appendix B for the formalism). To estimate its effect, another fit is performed with $A_{h}^{173}$, $B_{h}^{173}$, $\eta$, and $\zeta$. The result is summarized in the right column of Table 4. The A and B constants are consistent with the fit only with the A and B constants within a standard deviation inflated by $\chi^{2}/{\rm ndf}$. The amount of the second-order hyperfine interaction contributing to $A_{h}^{173}$, $B_{h}^{173}$, and $C_{h}^{173}$ is $5.1\pm 1.0$, $-22\pm 15$, and $-7.6\pm 1.8$ kHz, respectively. At the present accuracy of the spectroscopy, these numbers are comparable to the uncertainty with the inflation by $\sqrt{\chi^{2}/{\rm ndf}}$. Further improvement in the accuracy of the spectroscopy and inputs from theoretical calculations is desired to precisely determine the amount of second-order hyperfine interaction and whether the nucleus has a finite octupole moment. Note that same $\chi^{2}$ for the second and third column in Table 4 is not a coincidence. This is because the dependence of the second-order hyperfine interaction when $\zeta=-3\eta$ is the same as a linear combination of $\nu_{\rm ave}$, $A^{173}_{h}$, $B^{173}_{h}$, and $C^{173}_{h}$. The second-order hyperfine interaction shifts the center-of-mass frequency by $\-3.7(5.0)$ kHz, which is smaller than the uncertainty listed in Table 2. This is not large enough to explain the residual of the datapoint for 173Yb from the even-mass isotopes’ linearity in the King plot (See Sec. IV.3). It should be noted that in Ref. [30], analysis for $6s6p~{}^{3}P$ states based on a theoretical calculation of the wave function gives $\sim-500$-kHz shift for the center-of-mass frequency. It is possible that with such calculations, larger shift is predicted for the 431-nm transition as well. Further investigation on the effect of the second-order hyperfine structure on the shift in the center-of-mass frequency is necessary. In the following analysis, to be conservative, $A_{h}^{173}$ and $B_{h}^{173}$ obtained by the fit without $C^{173}_{h}$, $\eta$ and $\zeta$ are used.
Table 4: Hyperfine constants for odd-mass Yb isotopes in the unit of kHz: the top row specifies parameters included in the fit in each column. The uncertainty for 173Yb does not include the inflation by $\sqrt{\chi^{2}/{\rm ndf}}$. $A^{171}_{h}$ is cited from Ref. [22].
$A_{h},~{}B_{h}$ $A_{h},~{}B_{h},~{}C_{h}$ $A_{h},~{}B_{h},~{}\eta,~{}\zeta$
171Yb $A^{171}_{h}$ 1 123 354(13)
173Yb $A^{173}_{h}$309 449.1(2.0) -309 446.1(2.1) -309 441.0(2.6)
$B^{173}_{h}$1 700 631(22) -1 700 652(23) -1 700 674(27)
@@
$\eta^{173}_{h}$ 1 129(339)
$\zeta^{173}_{h}$3 774(895)
$\chi^{2}/{\rm ndf}$ 24.1/2 6.711/1 6.711/0
The ratio of $A_{h}$ between different isotopes equals the ratio of the nuclear $g$ factors $g^{A}$, as long as a pointlike nuclear dipole moment is assumed. The deviation $\Delta_{A,A^{\prime}}$ defined by the following equation is referred to as hyperfine anomaly [31]:
$\Delta_{171,173}=\frac{A_{h}^{171}/A_{h}^{173}}{g^{173}/g^{171}}-1=\frac{A_{h}^{171}/A_{h}^{173}}{(\mu_{n}^{173}/I^{173})/(\mu_{n}^{171}/I^{171})}-1,$ (5)
where $\mu_{n}^{A}$ and $I^{A}$ are the nuclear magnetic moment and nuclear spin, and $(\mu_{n}^{171}/I^{171})/(\mu_{n}^{173}/I^{173})=-3.6305(1)$ [32, 29, 33]. The hyperfine anomalies for the 431-nm and other transitions are summarized in Table 5. $\Delta_{171,173}$ for the $4f^{13}5d6s^{2}(J=2)$ state corresponding to the 431-nm transition is finite but particularly small compared to other states in Yb, as well as anomalies in most alkali-metal atoms [31]. Also, it is worth noting that the singlet states tend to have large anomaly.
Table 5: Hyperfine anomalies $\Delta_{171,173}$ for different states in Yb: references provide values for the A and B constants.
State $\Delta_{171,173}/10^{-3}$ Ref.
$6s6p~{}^{1}P_{1}$ 13.90(56) [34]
$4f^{13}5d6s^{2}(J=2)$0.084(55) this work
$6s6p~{}^{3}P_{2}$8.96(53) [35]
$6s6p~{}^{3}P_{1}$3.856(70) [36]
$6s7s~{}^{3}S_{1}$3.50(34) [35]
$6s5d~{}^{1}D_{2}$13.0(6.7) [37]
@@ IV.2 Nuclear charge radius
Table 6: Summary of $\lambda^{AA^{\prime}}$ and $\delta\langle r^{2}\rangle^{AA^{\prime}}$ with $A^{\prime}=176$: three digit numbers specify the transition used to obtain $\lambda^{AA^{\prime}}$ or $\delta\langle r^{2}\rangle^{AA^{\prime}}$. The columns without references are based on our data, except that the 578 isotope-shift data are cited from Ref. [20].
$A$ $\lambda^{AA^{\prime}}$ (fm2) $\delta\langle r^{2}\rangle^{AA^{\prime}}$ (fm2)
431 x-ray [38] 431 578 556 [30] Ref. [2]
168 -0.475 -0.505 -0.517 -0.4548(21) -0.550(128)
170 -0.336 -0.407(26) -0.357 -0.364 -0.3200(16) -0.392(38)
171 -0.287 -0.330(23) -0.305 -0.312 -0.2744(13) -0.329(77)
172 -0.205 -0.244(18) -0.218 -0.220 -0.1934(10) -0.233(24)
173 -0.158 -0.194(35) -0.168 -0.170 -0.1493(8) -0.180(24)
174 -0.100 -0.103(12) -0.107 -0.108 -0.0944(5) -0.117(30)
With theoretically calculated $F_{\alpha}$ and $K_{\alpha}$, differences in nuclear charge radii can be estimated using only the first two terms in Eq. (1). In this case where all higher order effects of nuclear charge are renormalized in $\delta\langle r^{2}\rangle^{AA^{\prime}}$, the parameter is referred to as the nuclear charge parameter, and denoted as $\lambda^{AA^{\prime}}$. The obtained numbers based on the 431-nm transition data are shown in Table 6. The relative uncertainty of $\delta\langle r^{2}\rangle^{AA^{\prime}}$ caused by the uncertainty of the spectroscopy is $\sim 10^{-5}$, and the uncertainty for the obtained $\delta\langle r^{2}\rangle^{AA^{\prime}}$ mainly originates from that of the theoretical calculation of $F_{\alpha}$ and $K_{\alpha}$, the amount of which is difficult to accurately estimate and therefore not shown in the table. Note that a conservative estimate discussed in Section III is ±$\pm 20$%.
To extract $\delta\langle r^{2}\rangle^{AA^{\prime}}$ from $\lambda^{AA^{\prime}}$, a relation $\delta\langle r^{2}\rangle^{AA^{\prime}}=\lambda^{AA^{\prime}}/0.941$ [30, 2] is utilized. The obtained $\delta\langle r^{2}\rangle^{AA^{\prime}}$ are shown in Table 6 within the uncertainty of the combined analysis in Ref. [2], but they tend to be smaller. One way to explain this is the discrepancy in $F_{431}/F_{578}$ between the theory and experiment. Empirically correcting $F_{431}$ to match the theoretical value of $F_{431}/F_{578}$ to the experimental value increases $\delta\langle r^{2}\rangle^{AA^{\prime}}$ by 13%. This reduces the discrepancy to Ref. [2], assuming that $F_{578}$ is reliable. Because $F_{578}$ and $K_{578}$ have a reasonable agreement between other reports, this assumption is reasonable to some extent, but the resulting $\delta\langle r^{2}\rangle^{AA^{\prime}}$ calculated by $F_{578}$ and $K_{578}$ obtained by our theory values and isotope shift in Ref. [20] shown in Table 6 is also smaller than Ref. [2].
In fact, such smaller $\delta\langle r^{2}\rangle^{AA^{\prime}}$ derived from optical spectroscopies compared to combined analysis like Refs. [2, 1] is observed in other reports as well [30, 29, 38]. Values obtained in Ref. [30], which is relatively new and has a careful assessment of their theoretical calculation within their own system, is shown in Table 6 as an example. The discrepancy seems to be because the x-ray spectroscopy provides relatively large $\lambda^{AA^{\prime}}$, as shown in Table 6 [38]. Spectroscopy of muonic atoms, which is another major contributor to the combined analysis, gives a close number to the calculation by the 431-nm transition data [38]. Some scattering experiments also favor $|\delta\langle r^{2}\rangle^{174,176}|$ smaller than 0.1 [39, 40, 41, 42]. For further interpretation of the data, a combined analysis with recent data is desired.
Figure 2: $\delta\langle r^{2}\rangle^{AA^{\prime}}$ calculated from isotope shifts with $A^{\prime}=176$: red circles show the values obtained from the absolute frequencies of the 431 nm transition, and black triangles show the values reported in Ref. [2]. The inset shows $\Delta\langle r^{2}\rangle^{(3)}$.
From these data, three-point odd-even staggering $\Delta\langle r^{2}\rangle^{(3)}$ is calculated as
@@
This is plotted in the inset of Fig. 2. Compared to the odd-even staggering observed in other atoms [43, 44, 45, 46], this effect is relatively large for stable isotopes. This could be attributed to the fact that both the number of protons and neutrons are far from the magic numbers, 50 and 82.
@@ IV.3 King Plot
Figure 3: Normalized King plot for the 431-nm transition: the plot is normalized with the 578-nm transition. Text near each data point shows the pair of isotopes that generated the data point. The blue line shows the linear fit of the black data points. See Table 3 for the fit parameters. The bottom half displays the residual $d^{A}$ of each data point from the fitted line. The residuals for the even-mass isotopes (black points) are magnified by a factor of 10.
The normalized King plot between the 431-nm transition and the 578-nm transition following Eq. 3 is shown in Fig. 3. In addition to the data points for even-mass isotopes, data points for odd-mass isotopes are plotted. The representative transition frequency for an odd-mass isotope is calculated as the center of gravity frequency of the hyperfine levels, and the isotope shift is calculated as the shift from the transition frequency for 174Yb, so that the data points lie between those of even-mass isotopes, and because 174Yb has smaller uncertainties compared to other even-mass isotopes. To assess the level of nonlinearity of the King plot, linear fit for the even-mass isotopes is performed by Eq. 3 with only the first two terms on the right-hand side. The plot exhibits slight nonlinearity beyond the standard deviation for some data points, as characterized by $\chi^{2}/{\rm ndf}=1.264$. However, as the probability for this level of deviation from $\chi^{2}/{\rm ndf}=1$ is 28.24%, the data points are linear within 95% confidence level.
@@
Although the amount of the nonlinearity is consistent with zero, the residual $d^{A}$ for isotope $A$ can be decomposed onto the two-dimensional space made by $\zeta_{\pm}=d^{168}-d^{170}\pm(d^{172}-d^{174})$, as introduced in Ref. [9]. The nonlinearity of the 431-nm transition shown in Fig. 4 is consistent with the one-parameter fit for the 361-, 411-, 436- and 467-nm transitions within one standard deviation. This finding supports the argument made in Ref. [3] that the major part of the nonlinearities in the King plots for Yb originates from several nuclear effects. This consistency observed between the previous results, as well as the almost linear behavior on the King plot, ensures that the measurements are not significantly affected by any systematic shifts beyond the listed uncertainties.
Table 7: Residual $\delta\nu^{A}$ from the linear fit of the King plot for odd-mass isotopes: the residual on the plot is multiplied by $\nu^{AA^{\prime}}_{578}$. $\lambda$ is the wavelength of the transition.
$\lambda$ (nm) $\delta\nu^{171}$ (kHz) $\delta\nu^{173}$ (kHz)
@@
$d^{A}$ for the odd-mass isotopes is significantly larger than the uncertainties on it, as shown in the bottom half of Fig. 3. To investigate the origin of this offset, the normalized King plots are made for the 399- and 556-nm transitions as well, and $\delta\nu^{A}=d^{A}\nu_{578}^{AA^{\prime}}$ is extracted, as shown in Table 7. On one hand, it is at most $\sim 1$ MHz and thus not significant in older experiments with large uncertainties [37, 47, 48]. On the other hand, with an accuracy of $\sim 10$ kHz, this is significant. $\delta\nu^{171}$, $\delta\nu^{173}$, and $\delta\nu^{173}/\delta\nu^{171}$ are plotted against $\Delta_{171,173}/10^{-3}$, $A_{h}^{171}$, $B_{h}^{173}$, $B_{h}^{173}/A_{h}^{173}$, and $F_{\alpha}$. Quantities with largest correlation are $B^{173}$ for $\delta\nu_{173}$ and $F_{\alpha}$ for $\delta\nu_{171}$, respectively. $\delta\nu^{173}/\delta\nu^{173}$ does not appear to have any good correlation against other quantities. However, it is not evident enough to state these quantities are related, because the $\chi^{2}/{\rm ndf}$ for the linear fit are 7.329 and 5.634, respectively, and both fitted lines do not cross the origin on the two-dimensional plot. One reason why the analysis is inconclusive is because $\delta\nu^{171}$ for the 399-nm transition has a large relative uncertainty. More precise spectroscopy, not only limited to the 399-, 431-, and 556-nm transitions but also for other narrow-linewidth transitions, is desired to figure out the source of the large $d^{A}$ for odd-mass isotopes.